# FVQ: A Large-Scale Dataset and A LMM-based Method for Face Video Quality Assessment

### v1

@@ Abstract.
Face video quality assessment (FVQA) deserves to be explored in addition to general video quality assessment (VQA), as face videos are the primary content on social media platforms and human visual system (HVS) is particularly sensitive to human faces. However, FVQA is rarely explored due to the lack of large-scale FVQA datasets. To fill this gap, we present the first large-scale in-the-wild FVQA dataset, FVQ-20K, which contains 20,000 in-the-wild face videos together with corresponding mean opinion score (MOS) annotations. Along with the FVQ-20K dataset, we further propose a specialized FVQA method named FVQ-Rater to achieve human-like rating and scoring for face video, which is the first attempt to explore the potential of large multimodal models (LMMs) for the FVQA task. Concretely, we elaborately extract multi-dimensional features including spatial features, temporal features, and face-specific features (i.e., portrait features and face embeddings) to provide comprehensive visual information, and take advantage of the LoRA-based instruction tuning technique to achieve quality-specific fine-tuning, which shows superior performance on both FVQ-20K and CFVQA datasets. Extensive experiments and comprehensive analysis demonstrate the significant potential of the FVQ-20K dataset and FVQ-Rater method in promoting the development of FVQA. The dataset will be released upon publication at: https://github.com/wsj-sjtu/FVQ.
Face video quality assessment, large multimodal model, dataset and benchmark
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06copyright: noneccs: Computing methodologies Computer visionccs: Human-centered computing Visualization design and evaluation methods
Figure 1. We explore the in-the-wild face video quality assessment problem for the first time. Concretely, we present FVQ-20K, the first large-scale in-the-wild face video quality assessment dataset, which contains 20,000 face videos with (a) diverse source video content, (b) various distortions in both spatial and temporal domains, (c) a variety of facial attributes, and (d) high-quality MOS annotation for each video. Along with the FVQ-20K dataset, we propose FVQ-Rater, the first LMM-based method elaborately designed for the face video quality assessment task as illustrated in (e).
@@ 1. Introduction
With the rapid development of network and video processing technologies, recent years have witnessed an explosion of social media applications with countless video content. However, these videos may suffer from various degradations caused by diverse acquisition and transmission conditions, thus video quality assessment (VQA) is becoming an increasingly popular topic in computer vision and multimedia with numerous research works (Ying et al., 2021; Lu et al., 2024a; Duan et al., 2024). Compared to general video content, face and portrait videos play an even more significant role in the quality of experience (QoE), since the human visual system (HVS) is particularly sensitive to human faces and has specialized mechanisms for face processing. Analyzing and measuring the quality of face videos can help monitor and optimize user experience at social media sites (Ying et al., 2021; Duan et al., 2024), promote the development of face-processing algorithms such as face video restoration, super-resolution, and enhancement, and ensure high-quality training data for generative models (Su et al., 2023; Chen et al., 2024a). However, face video quality assessment (FVQA) is rarely explored due to the absence of specific large-scale face video quality assessment datasets.
@@
Years of research on video quality assessment (VQA) have demonstrated promising results (You, 2021; Chen et al., 2021; Liao et al., 2022; Wu et al., 2022; Sun et al., 2022; Wu et al., 2023a, b; Yuan et al., 2023; Sun et al., 2024) thanks to the availability of various VQA datasets (Hosu et al., 2017; Sinno and Bovik, 2018; Wang et al., 2019; Ying et al., 2021; Lu et al., 2024a; Duan et al., 2024). However, these VQA datasets are designed for general videos, making the models devised on them less effective for face videos due to two aspects: (1) human perception of face video quality may differ from general video, as the HVS has specialized mechanisms for face processing (Kanwisher et al., 1997; Haxby et al., 2000; Su et al., 2023), which is overlooked in general VQA methods; (2) face videos only account for a small fraction of general VQA datasets, which limits the generalization ability of VQA methods to face videos. In recent years, some works have explored the problem of face image quality assessment from the aspects of either face recognition (Best-Rowden and Jain, 2018; Hernandez-Ortega et al., 2019; Terhorst et al., 2020; Ou et al., 2021; Boutros et al., 2023) or perceptual quality (Su et al., 2023; Chen et al., 2024a; Chahine et al., 2023; Liu et al., 2024b). However, they focus solely on face images without considering temporal distortions, making them less effective when directly applied to face videos. Recently, CFVQA (Li et al., 2024) has made the first attempt to explore the compressed FVQA problem. However, the dataset only contains 3,240 videos manually compressed from only 135 source videos using six video compression algorithms, which is not only limited in quantity and diversity but also fails to represent complex degradations in real-world scenarios. These limitations severely hinder the development of general FVQA algorithms.
@@
To alleviate the absence of general face video quality assessment datasets, we present FVQ-20K, the first large-scale in-the-wild face video quality assessment dataset (see Figure 1 and Table 1). Concretely, we collect face videos from both short-form video platform TikTok (Inc., 2016) and traditional video platform YouTube (YouTube, 2005) as shown in Figure 1 (a). For TikTok videos, we first collect over 100,000 videos across 11 categories related to people, and then filter and crop them to obtain 9,930 face videos, each with a fixed duration of 5 seconds. For YouTube videos, we integrate two public face video datasets, i.e., CelebV-HQ (Zhu et al., 2022) and CelebV-Text (Yu et al., 2023), and obtain a total of 10,070 face videos by filtering out duplicate identities and videos shorter than 5 seconds. After these steps, we obtain a rich in-the-wild face video dataset with various degradations such as blur, shaking, etc., as shown in Figure 1 (b), and diverse face attributes (Serengil and Ozpinar, 2024; Narayan et al., 2024) including gender, age, race, and emotion as shown in Figure 1 (c). Furthermore, we recruit 35 subjects from the university to rate the videos, and a professional team of image processing researchers is responsible for volunteer training and monitoring the quality of the scores. As shown in Figure 1 (d), the FVQ-20K dataset is equipped with high-quality mean opinion scores (MOSs) to represent the perceptual quality of each face video.
@@
  • We establish the first large-scale face video quality assessment dataset, FVQ-20K, which contains 20,000 diverse face videos with MOS annotations.
  • Based on the FVQ-20K dataset, we investigate and benchmark the performance of general-purpose VQA methods, FIQA methods, and state-of-the-art LMM models on the face video quality assessment task.
@@
2. Related Work
Face Quality Assessment. Face quality assessment contains face image quality assessment (FIQA) and face video quality assessment (FVQA). FIQA is an increasingly popular topic which can be divided into two research areas (Chen et al., 2024a; Su et al., 2023; Wang et al., 2022), including biometric face image quality assessment (BFIQA) (Abdel-Mottaleb and Mahoor, 2007; Gao et al., 2007; Abaza et al., 2012; Best-Rowden and Jain, 2018; Zhao et al., 2019; Lijun et al., 2019; Hernandez-Ortega et al., 2019; Terhorst et al., 2020; Meng et al., 2021; Ou et al., 2021; Boutros et al., 2023) and generic face image quality assessment (GFIQA) (Su et al., 2023; Chen et al., 2024a; Chahine et al., 2023; Liu et al., 2024b; Jo et al., 2023). BFIQA aims at ensuring the quality of face images for the face recognition systems (Schlett et al., 2022). Earlier BFIQA methods (Abdel-Mottaleb and Mahoor, 2007; Gao et al., 2007; Abaza et al., 2012; Lijun et al., 2019) usually evaluate the face image quality through hand-crafted features such as illumination, pose, facial expression, blur, occlusion, etc. Recent deep learning-based methods measure the quality of face images in a data-driven manner, utilizing either pre-trained face recognition backbones (Terhorst et al., 2020; Meng et al., 2021; Boutros et al., 2023) or human annotations (Best-Rowden and Jain, 2018; Zhao et al., 2019). However, these methods do not focus on the perceptual quality of face images. In contrast, GFIQA methods assess face images from the perspective of human perceptual quality, which is an emerging task (Su et al., 2023). Recently, impressive progress has been made in GFIQA thanks to the availability of several face image quality assessment datasets and advances in deep learning-based methods (Su et al., 2023; Chen et al., 2024a; Chahine et al., 2023; Liu et al., 2024b). Despite the extensive progress in FIQA, FVQA remains largely underexplored. The only work focusing on FVQA explores the perceptual quality in face video compression scenarios (Li et al., 2024) by collecting a compressed face video assessment dataset termed CFVQA. However, CFVQA only contains 3,240 videos manually compressed from 135 source videos using six video compression algorithms, which has limited quantity, diversity, and degradation types. Moreover, the method FAVOR devised based on CFVQA (Li et al., 2024) is a full-reference (FR) VQA method, which is hard to be used in real-world applications. To fill this gap, we construct the first large-scale in-the-wild FVQA dataset, FVQ-20K, and propose the first no-reference (NR) FVQA method, FVQ-Rater.
Video Quality Assessment. Video quality assessment (VQA) is a long-standing problem with extensive research efforts and a variety of publicly available datasets. VQA datasets can be categorized into three types according to the source of distortions in the videos, which includes captured VQA datasets (Nuutinen et al., 2016; Ghadiyaram et al., 2017), UGC VQA datasets (in-the-wild VQA datasets) (Hosu et al., 2017; Sinno and Bovik, 2018; Wang et al., 2019; Ying et al., 2021; Duan et al., 2024), and UGC+compression datasets (Li et al., 2020; Wang et al., 2021; Yu et al., 2021; Zhang et al., 2023; Lu et al., 2024a). However, all these datasets are designed for general-purpose videos and include very few face videos. VQA methods can be divided into two categories, including traditional methods (Saad et al., 2014; Mittal et al., 2015; Korhonen, 2019; Tu et al., 2021b; Kancharla and Channappayya, 2021; Tu et al., 2021a; Zheng et al., 2022) and learning-based methods (You, 2021; Chen et al., 2021; Liao et al., 2022; Wu et al., 2022; Sun et al., 2022; Wu et al., 2023a, b; Yuan et al., 2023; Sun et al., 2024). Traditional methods generally use hand-crafted features and predict quality scores through regression algorithms such as support vector regression (SVR) (Smola and Schölkopf, 2004), which struggle to handle complex real-world distortions. In recent years, learning-based methods have achieved superior performance through a data-driven manner. However, all these methods focus on general videos and do not consider human-specific features. In contrast, the proposed FVQ-20K dataset and the FVQ-Rater method are both specialized for face videos.
@@ 3. FVQ-20K Dataset
To address the scarcity of datasets and advance the progress of FVQA, we present FVQ-20K, the first large-scale in-the-wild face video quality assessment dataset as shown in Table 1, which contains 20,000 diverse face videos with high-quality MOS annotations. In this section, we introduce the dataset construction pipeline including face video collection (Section A.1), subjective study (Section 3.2), and subjective data processing (Section 3.3). The comprehensive analysis of the FVQ-20K dataset is detailed in Section 3.4.
@@ 3.3. Subjective Data Processing
where $r_{ij}$ is the raw score rated by the $i$-th subject to the $j$-th face video, $\mu_{i}$ and $\sigma_{i}$ represent the mean score and standard deviation given by subject $i$, respectively. Subsequently, Z-scores are linearly scaled to the range of $[0,100]$, assuming that the Z-scores of a subject follow a standard Gaussian distribution (Seshadrinathan et al., 2010; Li et al., 2024) within the range of $[-3,+3]$:
@@ 4. FVQ-Rater Method
, $\diamondsuit$, and $\clubsuit$ denote traditional VQA methods, learning-based VQA methods, BFIQA method, GFIQA method, LMM-based methods, and FVQA method, respectively. * and †indicate models trained and fine-tuned on the corresponding datasets, respectively. Note that FAVOR (Li et al., 2024) is a full-reference method and is not applicable to the in-the-wild FVQ-20K dataset. The best and runner-up performances are bold and underlined, respectively.
@@ 4.2. Training
Method / Dataset FVQ-20K (TikTok) FVQ-20K (YouTube) FVQ-20K CFVQA
@@ 5.1. Implementation Details
Evaluation Datasets. To the best of our knowledge, there are only two face video quality assessment datasets: CFVQA (Li et al., 2024) and our newly proposed FVQ-20K. We use both datasets in our experiments. Both datasets are randomly split into training, validation, and test sets with a ratio of $80\%:5\%:15\%$. Concretely, CFVQA dataset contains 3,240 videos, which is divided into 2,592 for training, 161 for validation, and 487 for testing. FVQ-20K contains 20,000 videos, with 16,000 for training, 1,000 for validation, and 3,000 for testing.
@@
Baseline Methods. Since there is no method specialized for the face video quality assessment (FVQA) task, we investigate and benchmark the state-of-the-art methods of three areas including FVQ, FIQA, and LMM on the FVQA task. The performance of the proposed FVQ-Rater is also evaluated by comparing with these methods: (1) general-purpose VQA methods: RAPIQUE (Tu et al., 2021b), VIDEVAL (Tu et al., 2021a), VSFA (Li et al., 2019), GSTVQA (Chen et al., 2021), SimpleVQA (Sun et al., 2022), FastVQA (Wu et al., 2022), Dover (Wu et al., 2023a), and KSVQE (Lu et al., 2024a); (2) FIQA methods: SDD-FIQA (Ou et al., 2021) and DSL-FIQA (Chen et al., 2024a); (3) LMM-based methods: Video-LLaVA (Lin et al., 2023), VideoLLaMA2 (Cheng et al., 2024), VideoLLaMA3 (Zhang et al., 2025), Qwen2-VL (Wang et al., 2024a), Qwen2.5-VL (Bai et al., 2025), Deepseek-VL (Lu et al., 2024b), InternVL2.5 (Chen et al., 2024b), and Q-Align (Wu et al., 2023c).

### v2

# FVQ: A Large-Scale Dataset and an LMM-based Method for Face Video Quality Assessment

@@ Abstract.
Face video quality assessment (FVQA) deserves to be explored in addition to general video quality assessment (VQA), as face videos are the primary content on social media platforms and human visual system (HVS) is particularly sensitive to human faces. However, FVQA is rarely explored due to the lack of large-scale FVQA datasets. To fill this gap, we present the first large-scale in-the-wild FVQA dataset, FVQ-20K, which contains 20,000 in-the-wild face videos together with corresponding mean opinion score (MOS) annotations. Along with the FVQ-20K dataset, we further propose a specialized FVQA method named FVQ-Rater to achieve human-like rating and scoring for face video, which is the first attempt to explore the potential of large multimodal models (LMMs) for the FVQA task. Concretely, we elaborately extract multi-dimensional features including spatial features, temporal features, and face-specific features (i.e., portrait features and face embeddings) to provide comprehensive visual information, and take advantage of the LoRA-based instruction tuning technique to achieve quality-specific fine-tuning, which shows superior performance on both FVQ-20K and CFVQA datasets. Extensive experiments and comprehensive analysis demonstrate the significant potential of the FVQ-20K dataset and FVQ-Rater method in promoting the development of FVQA. The code and dataset will be released at: https://github.com/wsj-sjtu/FVQ.
Face video quality assessment; large multimodal model; dataset and benchmark
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755332isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer visionccs: Human-centered computing Visualization design and evaluation methods
Figure 1. We explore the in-the-wild FVQA problem for the first time. Concretely, we present FVQ-20K, the first large-scale in-the-wild FVQA dataset, which contains 20,000 face videos with (a) diverse source video content, (b) various distortions in both spatial and temporal domains, (c) a variety of facial attributes, and (d) high-quality MOS annotation for each video. Along with the FVQ-20K dataset, we propose (e) FVQ-Rater, the first LMM-based method elaborately designed for the FVQA task.
@@ 1. Introduction
With the rapid development of network and video processing technologies, recent years have witnessed an explosion of social media applications with countless video content. However, these videos may suffer from various degradations caused by diverse acquisition and transmission conditions, thus video quality assessment (VQA) is becoming an increasingly popular topic in computer vision and multimedia with numerous research works (Ying et al., 2021; Lu et al., 2024a; Duan et al., 2024). Compared to general video content, face and portrait videos play an even more significant role in the quality of experience (QoE), as human faces naturally attract more visual attention than backgrounds or non-human objects, making their visual quality more perceptually significant. Analyzing and measuring the quality of face videos can help monitor and optimize user experience at social media sites (Ying et al., 2021; Duan et al., 2024), promote the development of face-processing algorithms such as face video restoration, super-resolution, and enhancement, and ensure high-quality training data for generative models (Su et al., 2023; Chen et al., 2024a). However, face video quality assessment (FVQA) is rarely explored due to the absence of specific large-scale FVQA datasets.
@@
Years of research on VQA have demonstrated promising results (You, 2021; Chen et al., 2021; Liao et al., 2022; Wu et al., 2022; Sun et al., 2022; Wu et al., 2023a, b; Yuan et al., 2023; Zhang et al., 2023a; Sun et al., 2024) thanks to the availability of various VQA datasets (Hosu et al., 2017; Sinno and Bovik, 2018; Wang et al., 2019; Ying et al., 2021; Lu et al., 2024a; Duan et al., 2024). However, these VQA datasets are designed for general videos, making the models devised on them less effective for face videos due to two aspects: (1) human perception of face video quality may differ from general video, as the HVS has specialized mechanisms for face processing (Kanwisher et al., 1997; Haxby et al., 2000; Su et al., 2023), which is overlooked in general VQA methods; (2) face videos only account for a small fraction of general VQA datasets, which limits the generalization ability of VQA methods to face videos. In recent years, some works have explored the problem of face image quality assessment from the aspects of either face recognition (Best-Rowden and Jain, 2018; Hernandez-Ortega et al., 2019; Terhorst et al., 2020; Ou et al., 2021; Boutros et al., 2023) or perceptual quality (Su et al., 2023; Chen et al., 2024a; Chahine et al., 2023; Liu et al., 2024b). However, they focus solely on face images without considering temporal distortions, making them less effective when directly applied to face videos. Recently, CFVQA (Li et al., 2024) has made the first attempt to explore the compressed FVQA problem. However, the dataset only contains 3,240 videos manually compressed from only 135 source videos using six video compression algorithms, which is not only limited in quantity and diversity but also fails to represent complex degradations in real-world scenarios. These limitations severely hinder the development of general FVQA algorithms.
@@
To alleviate the absence of general FVQA datasets, we present FVQ-20K, the first large-scale in-the-wild FVQA dataset (see Figure 1 and Table 1). Concretely, we collect face videos from both short-form video platform TikTok (Inc., 2016) and traditional video platform YouTube (YouTube, 2005) as shown in Figure 1 (a). For TikTok videos, we first collect over 100,000 videos across 11 categories related to people, and then filter and crop them to obtain 9,930 face videos, each with a fixed duration of 5 seconds. For YouTube videos, we integrate two public face video datasets, i.e., CelebV-HQ (Zhu et al., 2022) and CelebV-Text (Yu et al., 2023), and obtain a total of 10,070 face videos by filtering out duplicate identities and videos shorter than 5 seconds. After these steps, we obtain a rich in-the-wild face video dataset with various degradations such as blur, shaking, etc., as shown in Figure 1 (b), and diverse face attributes (Serengil and Ozpinar, 2024; Narayan et al., 2024) including gender, age, race, and emotion as shown in Figure 1 (c). Furthermore, we recruit 35 subjects from the university to rate the videos, and a professional team of image processing researchers is responsible for volunteer training and monitoring the quality of the scores. As shown in Figure 1 (d), the FVQ-20K dataset is equipped with high-quality mean opinion scores (MOSs) to represent the perceptual quality of each face video.
@@
  • We establish the first large-scale FVQA dataset, FVQ-20K, which contains 20,000 diverse face videos with MOS annotations.
  • Based on the FVQ-20K dataset, we investigate and benchmark the performance of general-purpose VQA methods, FIQA methods, and state-of-the-art LMM models on the FVQA task.
@@
2. Related Work
Face Quality Assessment. Face quality assessment contains face image quality assessment (FIQA) and face video quality assessment (FVQA). FIQA is an increasingly popular topic which can be divided into two research areas (Chen et al., 2024a; Su et al., 2023; Wang et al., 2022), including biometric face image quality assessment (BFIQA) (Abdel-Mottaleb and Mahoor, 2007; Gao et al., 2007; Abaza et al., 2012; Best-Rowden and Jain, 2018; Zhao et al., 2019; Lijun et al., 2019; Hernandez-Ortega et al., 2019; Terhorst et al., 2020; Meng et al., 2021; Ou et al., 2021; Boutros et al., 2023) and generic face image quality assessment (GFIQA) (Su et al., 2023; Chen et al., 2024a; Chahine et al., 2023; Liu et al., 2024b; Jo et al., 2023). BFIQA aims at ensuring the quality of face images for the face recognition systems (Schlett et al., 2022; Ou et al., 2024). Earlier BFIQA methods (Abdel-Mottaleb and Mahoor, 2007; Gao et al., 2007; Abaza et al., 2012; Lijun et al., 2019) usually evaluate the face image quality through hand-crafted features such as illumination, pose, facial expression, blur, occlusion, etc. Recent deep learning-based methods measure the quality of face images in a data-driven manner, utilizing either pre-trained face recognition backbones (Terhorst et al., 2020; Meng et al., 2021; Boutros et al., 2023) or human annotations (Best-Rowden and Jain, 2018; Zhao et al., 2019). However, these methods do not focus on the perceptual quality of face images. In contrast, GFIQA methods assess face images from the perspective of human perceptual quality, which is an emerging task (Su et al., 2023). Recently, impressive progress has been made in GFIQA thanks to the availability of several face image quality assessment datasets and advances in deep learning-based methods (Su et al., 2023; Chen et al., 2024a; Chahine et al., 2023; Liu et al., 2024b). Despite the extensive progress in FIQA, FVQA remains largely underexplored. The only work focusing on FVQA explores the perceptual quality in face video compression scenarios (Li et al., 2024) by collecting a compressed face video assessment dataset termed CFVQA. However, CFVQA only contains 3,240 videos manually compressed from 135 source videos using six video compression algorithms, which has limited quantity, diversity, and degradation types. Moreover, the method FAVOR devised based on CFVQA (Li et al., 2024) is a full-reference (FR) VQA method, which is hard to be used in real-world applications. To fill this gap, we construct the first large-scale in-the-wild FVQA dataset, FVQ-20K, and propose the first no-reference (NR) FVQA method, FVQ-Rater.
Video Quality Assessment. Video quality assessment (VQA) is a long-standing problem with extensive research efforts and a variety of publicly available datasets. VQA datasets can be categorized into three types according to the source of distortions in the videos, which includes captured VQA datasets (Nuutinen et al., 2016; Ghadiyaram et al., 2017), UGC VQA datasets (in-the-wild VQA datasets) (Hosu et al., 2017; Sinno and Bovik, 2018; Wang et al., 2019; Ying et al., 2021; Duan et al., 2024), and UGC+compression datasets (Li et al., 2020; Wang et al., 2021; Yu et al., 2021; Zhang et al., 2023b; Lu et al., 2024a). However, all these datasets are designed for general-purpose videos and include very few face videos. VQA methods can be divided into two categories, including traditional methods (Saad et al., 2014; Mittal et al., 2015; Korhonen, 2019; Tu et al., 2021b; Kancharla and Channappayya, 2021; Tu et al., 2021a; Zheng et al., 2022) and learning-based methods (You, 2021; Chen et al., 2021; Liao et al., 2022; Wu et al., 2022; Sun et al., 2022; Wu et al., 2023a, b; Yuan et al., 2023; Sun et al., 2024; Xing et al., 2024; Ge et al., 2025). Traditional methods generally use hand-crafted features and predict quality scores through regression algorithms such as support vector regression (SVR) (Smola and Schölkopf, 2004), which struggle to handle complex real-world distortions. In recent years, learning-based methods have achieved superior performance through a data-driven manner. However, all these methods focus on general videos and do not consider human-specific features. In contrast, the proposed FVQ-20K dataset and the FVQ-Rater method are both specialized for face videos.
@@ 3. FVQ-20K Dataset
To address the scarcity of datasets and advance the progress of FVQA, we present FVQ-20K, the first large-scale in-the-wild FVQA dataset as shown in Table 1, which contains 20,000 diverse face videos with high-quality MOS annotations. In this section, we introduce the dataset construction pipeline including face video collection (Section A.1), subjective study (Section 3.2), and subjective data processing (Section 3.3). The comprehensive analysis of the FVQ-20K dataset is detailed in Section 3.4.
@@ 3.3. Subjective Data Processing
where $r_{ij}$ is the raw score rated by the $i$-th subject to the $j$-th face video, $\mu_{i}$ and $\sigma_{i}$ represent the mean score and standard deviation given by subject $i$, respectively. Subsequently, Z-scores are linearly scaled to the range of $[0,100]$, assuming that the Z-scores of a subject follow a standard Gaussian distribution (Seshadrinathan et al., 2010; Li et al., 2024) within the range of [3,+3][-3,+3]:
@@ 3.4. Data Analysis
, $\diamondsuit$, and $\clubsuit$ denote traditional VQA methods, learning-based VQA methods, FIQA methods, FBP methods, LMM-based methods, and FVQA method, respectively. * and †indicate models trained and fine-tuned on the corresponding datasets, respectively. Note that FAVOR (Li et al., 2024) is a full-reference method and is not applicable to the in-the-wild FVQ-20K dataset. The best and runner-up performances are bold and underlined, respectively.
@@
REX-INCEP (Bougourzi et al., 2022)*
0.9130 0.9171 0.7467 0.9019 0.9052 0.7270 0.9174 0.9212 0.7528 0.4154 0.4333 0.3143
2D_FAP (Liu et al., 2025)*
0.6448 0.6655 0.4591 0.6783 0.6786 0.4887 0.6653 0.6734 0.4765 0.4229 0.4303 0.3012
MEBeauty (Lebedeva et al., 2022)*
0.7945 0.7976 0.5976 0.7971 0.8053 0.6052 0.7954 0.8023 0.6015 0.4278 0.4299 0.3064
$\diamondsuit$ Video-LLaVA-7B (Lin et al., 2023) 0.1338 0.1463 0.0989 0.0667 0.0950 0.0494 0.0766 0.0968 0.0565 0.0783 0.0613 0.0583
@@ 4.2. Training
Method / Dataset
FVQ-20K
FVQ-20K
FVQ-20K CFVQA
@@ 5.1. Implementation Details
Evaluation Datasets. To the best of our knowledge, there are only two FVQA datasets: CFVQA (Li et al., 2024) and our newly proposed FVQ-20K. We use both datasets in our experiments. Both datasets are randomly split into training, validation, and test sets with a ratio of $80\%:5\%:15\%$. Concretely, CFVQA dataset contains 3,240 videos, which is divided into 2,592 for training, 161 for validation, and 487 for testing. FVQ-20K contains 20,000 videos, with 16,000 for training, 1,000 for validation, and 3,000 for testing.
@@
Baseline Methods. Due to the lack of methods specifically tailored for the FVQA task, we systematically benchmark state-of-the-art methods from VQA, FIQA, Facial Beauty Prediction (FBP), and LMMs on the FVQA task, and conduct a comprehensive comparison with our proposed FVQ-Rater.